OpenStudy (emogirl100):

Simplify: Negative three and one-half minus negative twelve and one-fourth. Negative fifteen and three-fourths Fifteen and three-fourths Negative eight and three-fourths Eight and three-fourth

OpenStudy (queenie):

- 3 1/2 - (-12 1/4) = (turn to improper fractions) - 7/2 - (-49/4) = (negative times a negative is a positive) -7/2 + 49/4 = (find common denominator, which is 4 - 14/4 + 49/4 = (now just add..only the numerators, denominator stays the same) 35/4 = 8 3/4

OpenStudy (triciaal):

another approach same answer separate the whole and the fractions to work with smaller numbers